What mass of steam at 100 degrees C must be added to 1.00 kg of ice to yiekd liquid water at 30 degrees C?

 

Grade:12th pass

1 Answers

Eshan
5 years ago
Let the mass of steam required bem.

Heat lost by steam to reach temperature30^{\circ}Cis equal to the heat gained by ice to reach that temperature.

\implies m(540)+m(1)(100-30)=(1)(80)+(1)(1)(30-0)
\implies m=\dfrac{11}{61}kg